OpenStudy (anonymous):

multiply, answer in a+bi form: (4-5i)^2

jimthompson5910 (jim_thompson5910):

(4-5i)^2 = (4-5i)(4-5i) (4-5i)^2 = 4*4+4*(-5i)+(-5i)(4)+(-5i)(-5i) (4-5i)^2 = 16-20i-20i+25i^2 (4-5i)^2 = 16-20i-20i+25(-1) (4-5i)^2 = 16-20i-20i-25 (4-5i)^2 = -9-40i